# Perturbation-Restrained Sequential Model Editing

### v1

@@ Abstract
Model editing is an emerging field that focuses on updating the knowledge embedded within large language models (LLMs) without extensive retraining. However, current model editing methods significantly compromise the general abilities of LLMs as the number of edits increases, and this trade-off poses a substantial challenge to the continual learning of LLMs. In this paper, we first theoretically analyze that the factor affecting the general abilities in sequential model editing lies in the condition number of the edited matrix. The condition number of a matrix represents its numerical sensitivity, and therefore can be used to indicate the extent to which the original knowledge associations stored in LLMs are perturbed after editing. Subsequently, statistical findings demonstrate that the value of this factor becomes larger as the number of edits increases, thereby exacerbating the deterioration of general abilities. To this end, a framework termed Perturbation Restraint on Upper bouNd for Editing (PRUNE) is proposed, which applies the condition number restraints in sequential editing. These restraints can lower the upper bound on perturbation to edited models, thus preserving the general abilities. Systematically, we conduct experiments employing three popular editing methods on three LLMs across four representative downstream tasks. Evaluation results show that PRUNE can preserve considerable general abilities while maintaining the editing performance effectively in sequential model editing. The code and data are available at https://github.com/mjy1111/PRUNE.
@@ 1 Introduction
Despite the remarkable capabilities of large language models (LLMs), they encounter challenges such as false or outdated knowledge, and the risk of producing toxic content [65, 46, 29, 28]. Given the prohibitively high cost of retraining LLMs to address these issues, there has been a surge in focus on model editing [12, 41, 43, 44, 42, 39, 64, 27, 40], which aims at updating the knowledge of LLMs cost-effectively. Existing model editing methods can be roughly classified into either parameter-modifying methods  [43, 41, 42] that directly modify a small subset of model parameters, or parameter-preserving methods [44, 63] that integrate additional modules without altering the model parameters. In this paper, we study the parameter-modifying editing methods.
Sequential model editing involves making successive edits to the same model over time to continuously update knowledge, as illustrated in Figure 1(a). Recent studies [21, 22, 37] indicate that parameter-modifying editing methods significantly compromise the general abilities of LLMs as the number of edits increases, such as summarization, question answering, and natural language inference. However, these studies neither provide a theoretical analysis of the bottleneck of the general abilities of the edited models, nor propose a solution to preserve these abilities in sequential editing. These affect the scalability of model editing and pose a substantial challenge to the continual learning of LLMs.
@@
In light of the above issues, we first theoretically analyze through matrix perturbation theory [38, 54, 60] to elucidate a crucial factor affecting the general abilities during sequential editing: the condition number [50, 13, 52] of the edited matrix. The condition number of a matrix represents its numerical sensitivity and therefore can be used to indicate the extent to which the original knowledge associations stored in LLMs are perturbed after editing. As shown in Figure 1(b), our statistical findings demonstrate that the condition number of the edited matrix substantially increases as the number of edits increases, thereby exacerbating the perturbation of original knowledge and the deterioration of general abilities. Therefore, we assume that the bottleneck of the general abilities during sequential editing lies in the escalating value of the condition number.
@@
Towards continual and scalable model editing, we propose Perturbation Restraint on Upper bouNd for Editing (PRUNE) based on the above analysis, which applies the condition number restraints in sequential editing to preserve general abilities and maintain new editing knowledge simultaneously. Specifically, the condition number of the edited matrix is restrained by reducing the large singular values [1, 57] of the edit update matrix. Consequently, the upper bound on perturbation to the edited matrix is lowered, thus reducing the perturbation to the original knowledge associations and preserving the general abilities of the edited model, as shown in Figure 1(c). Additionally, we observe that these larger singular values often encapsulate redundant editing overfitting information, so regularizing them will not affect the newly editing knowledge, as shown in Figure 1(d). In this way, the new editing knowledge is embedded into LLMs without affecting their original general abilities. Overall, the proposed editing framework requires only minimal computing resources, and is adaptable to be coupled with multiple existing editing methods.
@@
To validate the effectiveness of the proposed PRUNE, our study comprehensively evaluates the edited LLMs for both general abilities and editing performance in sequential editing scenarios. Extensive empirical research involves three popular editing methods, including MEND [43], ROME [41], and MEMIT [42], which are analyzed based on three representative LLMs including GPT-2 XL (1.5B) [48], LLaMA-2 (7B) [53], and LLaMA-3 (8B). Four representative downstream tasks including reasoning [8], summarization [19], open-domain QA [31], and natural language inference [11] are employed to extensively demonstrate the impact of model editing on the general abilities of LLMs. Experimental results demonstrate that the proposed PRUNE can preserve considerable general abilities and maintain almost all editing performance in sequential editing.
@@
In essence, our research offers three significant contributions: (1) This study theoretically analyzes that the escalating value of the condition number of the edited matrix is the bottleneck of sequential model editing. (2) The PRUNE framework based on the analysis is proposed to preserve the general abilities of the edited model while retaining the editing knowledge. (3) Experimental results including both editing performance and four downstream task performance across three editing methods on three LLMs demonstrate the effectiveness of the proposed method.
@@ 2 Related Work
From the perspective of whether the model parameters are modified, existing editing methods can be divided into parameter-modifying  [43, 41, 42, 12] and parameter-preserving methods [44, 25, 63]. This paper focuses on the former. Previous works have investigated the role of MLP layers in Transformer, showing that MLP layers store knowledge, which can be located in specific neurons and edited [16, 10, 17]. KE [3] and MEND [43] train a hypernetwork to get gradient changes to update model parameters [43]. Besides, Meng et al. [41] and Meng et al. [42] used Locate-Then-Edit strategy, which first located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair in the MLP module. Parameter-preserving methods do not modify model weights but store the editing facts with an external memory. For example, Mitchell et al. [44] stored edits in a base model and learned to reason over them to adjust its predictions as needed.
@@
Some works investigate the paradigm for model editing evaluation [67, 9, 39, 35, 26, 61, 15, 40]. Cohen et al. [9] introduced the ripple effects of model editing, suggesting that editing a particular fact implies that many other facts need to be updated. Ma et al. [39] constructed a new benchmark to assess the edited model bidirectionally. Besides, Li et al. [35] explored two significant areas of concern: Knowledge Conflict and Knowledge Distortion. These early studies mainly evaluate edited models per edit rather than sequentially, and they focus narrowly on basic factual triples. Recently, some works assess the impact of editing methods on the general abilities of LLMs in sequential editing scenarios. These studies [21, 22, 37] have conducted comprehensive experiments, showing the parameter-modifying methods significantly degrade the model performance on general downstream tasks.
@@
It plays a crucial role in the field of artificial intelligence (AI) by providing a systematic framework to understand the impact of small changes or perturbations in various AI algorithms and models. Some studies [24, 47, 49] delve into the interpretability of LLMs, revealing how minor alterations in input features or model parameters influence the model’s predictions. This understanding helps uncover significant feature connections within the model architecture. Moreover, it has been instrumental in assessing and enhancing the robustness of models [5, 20, 6]. Furthermore, Bird et al. [2] and Dettmers et al. [14] have employed it for sensitivity analysis to identify critical factors affecting algorithm performance. It also contributes to the development of efficient optimization techniques [34, 7, 30], improving convergence rates and stability of optimization algorithms.
Compared with previous works [41, 42, 62, 21, 22, 37] that are the most relevant, a main difference should be highlighted. They neither theoretically investigate the reasons for general ability degradation, nor propose methods to maintain these abilities during sequential editing. In contrast, our study makes the first attempt to theoretically explore the bottleneck of general abilities in sequential editing and proposes the PRUNE framework to preserve these abilities for continual model editing.
@@ 3.1 Preliminary
This task involves modifying the memorized knowledge contained in LMs. Various kinds of complex learned beliefs such as logical, spatial, or numerical knowledge are expected to be edited. In this paper, following previous work [41, 67, 42, 64], we study editing factual knowledge in the form of (subject s, relation r, object o), e.g., (s = United States, r = President of, o = Donald Trump). An LM is expected to recall a memory representing $o$ given a natural language prompt $p(s,r)$ such as “The President of the United States is”. Editing a fact is to incorporate a new knowledge triple $(s,r,o^{*})$ in place of the current one $(s,r,o)$. An edit is represented as $e=(s,r,o,o^{*})$ for brevity. Given a set of editing facts =$\mathcal{E}=\left\{e_{1},e_{2},\ldots\right\}$ and an original model $f_{\theta_{0}}$, sequential model editing operationalizes each edit after the last edit [^1] This paper studies editing a single fact at a time and leaves the exploration of batch editing as future work., i.e., $K(f_{\theta_{n-1}},e_{n})=f_{\theta_{n}}$, where $f_{\theta_{n}}$ denotes the model after $n$ edits.
@@
SVD [1] is a fundamental and effective matrix factorization technique for analyzing matrix structures. Formally, an SVD of a matrix $W\in\mathbb{R}^{p\times q}$ is given by $W=U\Sigma V^{\mathrm{T}}$, where $U=[u_{1},u_{2},...,u_{p}]\in\mathbb{R}^{p\times p}$, $V=[v_{1},v_{2},...,v_{q}]\in\mathbb{R}^{q\times q}$, and $\Sigma\in\mathbb{R}^{p\times q}$. $u_{i}$ and $v_{i}$ are the column vectors of $U$ and $V$, and constitute an orthonormal basis of $\mathbb{R}^{p}$ and $\mathbb{R}^{q}$ respectively. $\Sigma$ is a diagonal matrix whose diagonal entries are given by the singular values of $W$ in descending order. Additionally, the SVD of $W$ could also be formulated as: $W=\sum_{i=1}^{\min\{p,q\}}\sigma_{i}u_{i}v_{i}^{\mathrm{T}}$, where $\sigma_{i}$ is singular value, and $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{\min\{p,q\}}\geq 0$. In the scenario of this paper, $W$ is a full-rank matrix, so $\sigma_{\min\{p,q\}}>0$.
@@ 3.2 Matrix Perturbation Theory Analysis
Previous works [16, 41, 23, 59] have analyzed and located that the MLP modules in Transformer [55] store various kinds of knowledge [45, 56]. The MLP module of the $l$-th Transformer layer consists of two projection layers, where the first and second layers are denoted as $W_{fc}^{l}$ and $W_{proj}^{l}$ respectively. $W_{proj}^{l}$ is considered as a linear associative memory which stores knowledge in the form of key-value pairs $(k_{i},v_{i})$, and is usually regarded as the editing area [41, 42]. In this paper, $W_{proj}^{l}$ is denoted as $W$ for brevity. $W$ is assumed to store many key-value pairs $P=\{(k_{i},v_{i})\mid i=1,2,...\}$ which satisfies $Wk_{i}=v_{i}$, where $k_{i}\in\mathbb{R}^{q}$ and $v_{i}\in\mathbb{R}^{p}$. Assuming $\mid\mathcal{E}\mid=N$ in sequential model editing, an edit update matrix $\Delta W_{j}$ is calculated for the edit $e_{j}$ and added to $W$, which can be formulated as: $W_{N}=W+\sum_{j=1}^{N}\Delta W_{j}$ with $\Delta W_{j}$ calculated from $f_{\theta_{j-1}}$.
@@
3.3 Change Trend of Condition Number
@@
Surprisingly, we observed that regardless of the editing methods employed, the condition number of the edited matrix exhibited a rapid increase as the number of edits increased, especially after a large number of edits. According to Theorem 3.1, the adjustment norm $\|\Delta k^{n}_{i}\|_{2}$ corresponding to the $n$-th edit tends to increase as the number of edits $n$ increases. It underscores that when the number of edits is relatively large, the upper bound on the perturbation caused by subsequent edits to the key-value pair associations becomes very large, further disrupting the stored original knowledge and exacerbating the deterioration of general abilities.
Figure 2: The condition number, maximum singular value and minimum singular value of the edited matrix in sequential editing. Three editing methods including ROME, MEND, and MEMIT are used to edit LLaMA-2 (7B) on the CounterFact [41] dataset. For editing methods that modify the parameters of multiple MLP layers, one of them is randomly selected for illustration. $W$ and $W_{n}$ denote the unedited and edited matrices respectively.

### v3

@@ Abstract
Model editing is an emerging field that focuses on updating the knowledge embedded within large language models (LLMs) without extensive retraining. However, current model editing methods significantly compromise the general abilities of LLMs as the number of edits increases, and this trade-off poses a substantial challenge to the continual learning of LLMs. In this paper, we first theoretically analyze that the factor affecting the general abilities in sequential model editing lies in the condition number of the edited matrix. The condition number of a matrix represents its numerical sensitivity, and therefore can be used to indicate the extent to which the original knowledge associations stored in LLMs are perturbed after editing. Subsequently, statistical findings demonstrate that the value of this factor becomes larger as the number of edits increases, thereby exacerbating the deterioration of general abilities. To this end, a framework termed Perturbation Restraint on Upper bouNd for Editing (PRUNE) is proposed, which applies the condition number restraints in sequential editing. These restraints can lower the upper bound on perturbation to edited models, thus preserving the general abilities. Systematically, we conduct experiments employing three editing methods on three LLMs across four downstream tasks. The results show that PRUNE can preserve general abilities while maintaining the editing performance effectively in sequential model editing. The code are available at https://github.com/mjy1111/PRUNE.
@@ 1 Introduction
Despite the remarkable capabilities of large language models (LLMs), they encounter challenges such as false or outdated knowledge, and the risk of producing toxic content (Zhang et al., 2023; Peng et al., 2023; Ji et al., 2023; Huang et al., 2023). Given the prohibitively high cost of retraining LLMs to address these issues, there has been a surge in focus on model editing (Dai et al., 2022; Meng et al., 2022; Mitchell et al., 2022a; b; Meng et al., 2023; Zhang et al., 2024; Hu et al., 2024; Ma et al., ), which aims at updating the knowledge of LLMs cost-effectively. Existing model editing methods can be roughly classified into either parameter-modifying methods  (Mitchell et al., 2022a; Meng et al., 2022; 2023) that directly modify a small subset of model parameters, or parameter-preserving methods (Mitchell et al., 2022b; Yu et al., 2024) that integrate additional modules without altering the model parameters. In this paper, we study the parameter-modifying editing methods.
Sequential model editing involves making successive edits to the same model over time to continuously update knowledge, as illustrated in Figure 1(a). Recent studies (Gu et al., 2024; Gupta et al., 2024b; Lin et al., 2024; Gupta et al., 2024a) indicate that parameter-modifying editing methods significantly compromise the general abilities of LLMs as the number of edits increases, such as summarization, question answering, and natural language inference. However, these studies neither provide a theoretical analysis of the bottleneck of the general abilities of the edited models, nor propose a solution to preserve these abilities in sequential editing. These affect the scalability of model editing and pose a substantial challenge to the continual learning of LLMs.
@@
In light of the above issues, we first theoretically analyze through matrix perturbation theory (Vaccaro, 1994; Wedin, 1972) to elucidate a crucial factor affecting the general abilities during sequential editing: the condition number (Smith, 1967; Dedieu, 1997; Sun, 2000) of the edited matrix. The condition number of a matrix represents its numerical sensitivity and therefore can be used to indicate the extent to which the original knowledge associations stored in LLMs are perturbed after editing. As shown in Figure 1(b), statistical findings demonstrate that the condition number of the edited matrix substantially increases as the number of edits increases, thereby exacerbating the perturbation of original knowledge and the deterioration of general abilities. Therefore, we assume that the bottleneck of the general abilities during sequential editing lies in the escalating value of the condition number.
@@
Towards continual and scalable model editing, we propose Perturbation Restraint on Upper bouNd for Editing (PRUNE) based on the above analysis, which applies the condition number restraints in sequential editing to preserve general abilities and maintain new editing knowledge simultaneously. Specifically, the condition number of the edited matrix is restrained by reducing the large singular values (Albano et al., 1988; Wall et al., 2003) of the edit update matrix. Consequently, the upper bound on perturbation to the edited matrix is lowered, thus reducing the perturbation to the original knowledge associations and preserving the general abilities of the edited model, as shown in Figure 1(c). Additionally, we observe that these larger singular values often encapsulate redundant editing overfitting information, so regularizing them will not affect the newly editing knowledge, as shown in Figure 1(d). In this way, the new editing knowledge is embedded into LLMs without affecting their original general abilities. Overall, the proposed editing framework requires only minimal computing resources, and is adaptable to be coupled with multiple existing editing methods.
@@
To validate the effectiveness of the proposed PRUNE, our study comprehensively evaluates the edited LLMs for both general abilities and editing performance in sequential editing scenarios. Extensive research involves three popular editing methods, including MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023), which are analyzed based on three LLMs including GPT-2 XL (1.5B) (Radford et al., 2019), LLaMA-2 (7B) (Touvron et al., 2023), and LLaMA-3 (8B). Four downstream tasks including reasoning (Cobbe et al., 2021), summarization (Gliwa et al., 2019), open-domain QA (Kwiatkowski et al., 2019), and natural language inference (Dagan et al., 2005) are employed to demonstrate the impact of editing on the general abilities of LLMs. Experimental results demonstrate that the proposed PRUNE can preserve considerable general abilities and maintain almost all editing performance in sequential editing.
@@
In essence, our research offers three contributions: (1) This study theoretically analyzes that the escalating value of the condition number of the edited matrix is the bottleneck of sequential model editing. (2) The PRUNE based on the analysis is proposed to preserve the general abilities of the edited model while retaining the editing knowledge. (3) Experimental results including both editing and downstream task performance across three editing methods on three LLMs demonstrate the effectiveness of PRUNE.
@@ 2 Related Work
From the perspective of whether the model parameters are modified, existing editing methods can be divided into parameter-modifying  (Mitchell et al., 2022a; Meng et al., 2022; 2023; Dai et al., 2022) and parameter-preserving methods (Mitchell et al., 2022b; Hartvigsen et al., 2023; Yu et al., 2024). This paper focuses on the former. Previous works have investigated the role of MLP layers in Transformer, showing that MLP layers store knowledge, which can be located in specific neurons and edited (Geva et al., 2021; Da et al., 2021; Geva et al., 2022). KE (Cao et al., 2021) and MEND (Mitchell et al., 2022a) train a hypernetwork to get gradient changes to update model parameters (Mitchell et al., 2022a). Besides, Meng et al. (2022) and Meng et al. (2023) used Locate-Then-Edit strategy, which first located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair in the MLP module. Parameter-preserving methods do not modify model weights but store the editing facts with an external memory. For example, Mitchell et al. (2022b) stored edits in a base model and learned to reason over them to adjust its predictions as needed.
@@
Some works investigate the paradigm for model editing evaluation (Zhong et al., 2023; Cohen et al., 2024; Ma et al., 2023; Li et al., ; Hase et al., 2023; Wu et al., 2023; Gandikota et al., 2023; Ma et al., ). Cohen et al. (2024) introduced the ripple effects of model editing, suggesting that editing a particular fact implies that many other facts need to be updated. Ma et al. (2023) constructed a new benchmark to assess the edited model bidirectionally. Besides, Li et al. explored two significant areas of concern: Knowledge Conflict and Knowledge Distortion. These early studies mainly evaluate edited models per edit rather than sequentially, and they focus narrowly on basic factual triples. Recently, some works assess the impact of editing methods on the general abilities of LLMs in sequential editing scenarios. These studies (Gu et al., 2024; Gupta et al., 2024b; Lin et al., 2024; Yang et al., 2024; Gupta et al., 2024a; c) have conducted comprehensive experiments, showing the parameter-modifying methods significantly degrade the model performance on downstream tasks.
@@
It plays a crucial role in the field of artificial intelligence (AI) by providing a systematic framework to understand the impact of small changes or perturbations in various AI algorithms and models. Some studies (Harder et al., 2020; Qin et al., 2022; Singh et al., 2024) delve into the interpretability of LLMs, revealing how minor alterations in input features or model parameters influence the model’s predictions. This understanding helps uncover significant feature connections within the model architecture. Moreover, it has been instrumental in assessing and enhancing the robustness of models (Chen et al., 2023; 2024). Furthermore, Bird et al. (2020) and Dettmers et al. (2023) have employed it for sensitivity analysis to identify critical factors affecting algorithm performance. It also contributes to the development of efficient optimization techniques (Li et al., 2020; Jiang et al., 2024), improving convergence rates and stability of optimization algorithms.
Compared with previous works (Meng et al., 2022; 2023; Yao et al., 2023; Gu et al., 2024; Gupta et al., 2024b; Lin et al., 2024) that are the most relevant, a main difference should be highlighted. They neither theoretically investigate the reasons for general ability degradation, nor propose effective methods to maintain these abilities during sequential editing. In contrast, our study makes the first attempt to theoretically explore the bottleneck of general abilities in sequential editing and proposes the PRUNE framework to preserve these abilities for continual model editing.
@@ 3.1 Preliminary
This task involves modifying the memorized knowledge contained in LMs. Various kinds of complex learned beliefs such as logical, spatial, or numerical knowledge are expected to be edited. In this paper, following previous work (Meng et al., 2022; Zhong et al., 2023; Meng et al., 2023; Zhang et al., 2024), we study editing factual knowledge in the form of (subject s, relation r, object o), e.g., (s = United States, r = President of, o = Donald Trump). An LM is expected to recall a memory representing $o$ given a natural language prompt $p(s,r)$ such as “The President of the United States is”. Editing a fact is to incorporate a new knowledge triple $(s,r,o^{*})$ in place of the current one $(s,r,o)$. An edit is represented as $e=(s,r,o,o^{*})$ for brevity. Given a set of editing facts =$\mathcal{E}=\left\{e_{1},e_{2},\ldots\right\}$ and an original model $f_{\theta_{0}}$, sequential model editing operationalizes each edit after the last edit [^1] This paper studies editing a single fact at a time and leaves the exploration of batch editing as future work., i.e., $K(f_{\theta_{n-1}},e_{n})=f_{\theta_{n}}$, where $f_{\theta_{n}}$ denotes the model after $n$ edits.
@@
SVD (Albano et al., 1988) is a fundamental and effective matrix factorization technique for analyzing matrix structures. Formally, an SVD of a matrix $W\in\mathbb{R}^{p\times q}$ is given by $W=U\Sigma V^{\mathrm{T}}$, where $U=[u_{1},u_{2},...,u_{p}]\in\mathbb{R}^{p\times p}$, $V=[v_{1},v_{2},...,v_{q}]\in\mathbb{R}^{q\times q}$, and $\Sigma\in\mathbb{R}^{p\times q}$. $u_{i}$ and $v_{i}$ are the column vectors of $U$ and $V$, and constitute an orthonormal basis of $\mathbb{R}^{p}$ and $\mathbb{R}^{q}$ respectively. $\Sigma$ is a diagonal matrix whose diagonal entries are given by the singular values of $W$ in descending order. Additionally, the SVD of $W$ could also be formulated as: $W=\sum_{i=1}^{\min\{p,q\}}\sigma_{i}u_{i}v_{i}^{\mathrm{T}}$, where $\sigma_{i}$ is singular value, and $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{\min\{p,q\}}\geq 0$. In the scenario of this paper, $W$ is a full-rank matrix, so $\sigma_{\min\{p,q\}}>0$.
@@ 3.2 Matrix Perturbation Theory Analysis
Previous works (Geva et al., 2021; Meng et al., 2022; Gupta et al., 2023; Wang et al., 2024) have analyzed and located that the MLP modules in Transformer (Vaswani et al., 2017) store various kinds of knowledge (Pearl, 2001; Vig et al., 2020). The MLP module of the $l$-th Transformer layer consists of two projection layers, where the first and second layers are denoted as $W_{fc}^{l}$ and $W_{proj}^{l}$ respectively. $W_{proj}^{l}$ is considered as a linear associative memory which stores knowledge in the form of key-value pairs $(k_{i},v_{i})$, and is usually regarded as the editing area (Meng et al., 2022; 2023). In this paper, $W_{proj}^{l}$ is denoted as $W$ for brevity. $W$ is assumed to store many key-value pairs $P=\{(k_{i},v_{i})\mid i=1,2,...\}$ which satisfies $Wk_{i}=v_{i}$, where $k_{i}\in\mathbb{R}^{q}$ and $v_{i}\in\mathbb{R}^{p}$. Assuming $\mid\mathcal{E}\mid=N$ in sequential model editing, an edit update matrix $\Delta W_{j}$ is calculated for the edit $e_{j}$ and added to $W$, which can be formulated as: $W_{N}=W+\sum_{j=1}^{N}\Delta W_{j}$ with $\Delta W_{j}$ calculated from $f_{\theta_{j-1}}$.
@@
3.3 Trend of the Condition Number During Sequential Editing
@@
Surprisingly, we observed that regardless of the editing methods employed, the condition number of the edited matrix exhibited a rapid increase as the number of edits increased, particularly after a large number of edits. According to Theorem 3.1, the adjustment norm $\|\Delta k^{n}_{i}\|_{2}$ corresponding to the $n$-th edit tends to increase as the number of edits $n$ increases. Therefore, we can draw two conclusions: (1) As more edits are performed, the upper bound of the perturbation caused by a new single edit to the key-value pair associations increases. (2) During the sequential model editing process, the cumulative perturbation of these edits will become larger and larger. These factors further disrupt the stored original knowledge and exacerbate the deterioration of general abilities. As the second conclusion is easy to understand, here is an example for the first point. From the first subfigure of Figure 2, we can observe that the condition number of the $W_{200}$ matrix after the 200th edit is significantly higher than that of the unedited matrix $W$. Therefore, the perturbation of the model caused by the 201st edit is likely to be much greater than the perturbation of the model caused by the 1st edit.
Figure 2: The condition number, maximum singular value and minimum singular value of the edited matrix in sequential editing. Three editing methods including ROME, MEND, and MEMIT are used to edit LLaMA-2 (7B) on the CounterFact (Meng et al., 2022) dataset. For editing methods that modify the parameters of multiple MLP layers, one of them is randomly selected for illustration. $W$ and $W_{n}$ denote the unedited and edited matrices respectively.